# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v3

@@ 1 Introduction
Perception is critical in autonomous driving. While many single-vehicle perception methods [32; 31; 15; 27; 61; 47; 25; 24; 28; 14; 48; 30] have emerged, substantial safety challenges persist due to occlusions and limited perception ranges. These issues occur because vehicles can only view their surroundings from a single perspective, resulting in an incomplete understanding of the scenario. This limitation hinders autonomous vehicles from achieving safe navigation and making optimal decisions. To address this challenge, recent studies have explored cooperative perception, where ego vehicles extend their perception range and overcome occlusions with assistance from other vehicles or roadside perception [45; 44; 46; 49; 34].
@@
  • We have developed V2X-Radar, the first extensive real-world multi-modal dataset featuring 4D Radar designed for cooperative perception. All frames are captured through a variety of real-world scenarios using multi-modal sensors.
@@
  • Comprehensive benchmarks of recent perception algorithms are conducted across cooperative perception on V2X-Radar-C, roadside perception on V2X-Radar-I, and vehicle-side perception on V2X-Radar-V. We will release the entire V2X-Radar dataset and benchmark codebase.
@@ 2.1 Autonomous Driving Datasets
Public datasets [1; 29; 42; 41; 13; 56; 54; 40; 35; 19; 59; 60] have significantly accelerated the progress of autonomous driving in recent years. Single-vehicle datasets, such as KITTI [6], NuScenes [1], and Waymo [29], have notably furthered the development of single-vehicle perception. In terms of cooperative perception datasets, OPV2V [42] is the first dataset in this area, gathering data through co-simulation using CARLA [3] and OpenCDA [37; 38]. V2XSet [41] and V2X-Sim [13] explore V2X perception by employing synthesized data from the CARLA simulator [3]. Unlike simulated datasets, DAIR-V2X introduces the first real-world dataset for cooperative detection. V2X-Seq [56] extends sequences from DAIR-V2X [54] with track IDs, establishing a sequential perception and trajectory forecasting dataset. V2V4Real [40] presents the initial real-world V2V dataset gathered from two connected vehicles. V2X-Real [35] is a large-scale multi-modal multi-view dataset intended for V2X research, compiled using two connected automated vehicles and two intelligent roadside units. TUMTraf-V2X [62] is a multimodal, multi-view V2X cooperative perception dataset designed for 3D object detection and tracking in traffic scenarios. A common limitation among the cooperative perception datasets mentioned above is their narrow focus on camera and LiDAR sensors, overlooking the benefits of 4D Radar. 4D Radar is recognized for its superior adaptability to adverse weather conditions, as demonstrated by single-vehicle autonomous driving datasets such as K-Radar [19] and Dual-Radar [59]. Consequently, It’s essential to create a multimodal dataset that integrates 4D Radar to enhance cooperative perception research.
@@ 2.2 Cooperative Perception
Cooperative perception aims to extend the range of perception and overcome occlusions in single-vehicle perception by leveraging shared information among connected agents. It can be classified into three main types based on fusion strategies: (1) Early Fusion, which involves transmitting raw sensor data for the ego vehicle to aggregate and predict objects; (2) Late Fusion, which integrates detection results for a consistent prediction; and (3) Intermediate Fusion, which shares and fuses intermediate high-level features. Recent state-of-the-art methods [33; 2; 39; 41; 8; 17; 33; 4; 55; 57; 5] typically follow intermediate fusion, achieving the best trade-off between accuracy and bandwidth requirements. For instance, V2VNet [33] uses a graph neural network to refine features and perform joint perception and prediction iteratively. F-Cooper [2] introduces a pooling mechanism for identifying salient features. CoBEVT [39] introduced local-global sparse attention for improved cooperative BEV map segmentation. V2X-ViT [41] presented a unified vision transformer for multi-agent and multi-scale perception. HM-ViT [8] introduced a 3D heterogeneous graph transformer for camera and LiDAR fusion. FFNet [55] is a novel flow-based feature fusion framework that addresses temporal asynchrony in cooperative 3D object detection through feature flow prediction. HEAL [17] proposed an extensible collaborative perception framework to integrate new heterogeneous agents with minimal integration cost and performance decline.
@@ 4.2 Cooperative 3D Object Detection
Evaluation Metrics. The evaluation area spans [-100, 100] metres in both the x and y directions relative to the ego vehicle. Similar to DAIR-V2X [54] and V2V4Real [40], we group various vehicle types into the same class and concentrate solely on vehicle detection. The performance of object detection is evaluated using Average Precision (AP) at IoU thresholds of 0.5 and 0.7. Transmission costs are calculated through Average MegaBytes (AM). In line with previous studies [54; 40], we compare methods in two configurations: (1) Synchronous, ignoring communication delays. (2) Asynchronous, simulating the delay by retrieving roadside samples with the preceding timestamp.
@@ 5.2 Benchmark Results
Method M Vehicle (IoU = 0.7 / 0.5) Pedestrian (IoU = 0.5 / 0.25) Cyclist (IoU = 0.5 / 0.25)
@@
Method M Vehicle (IoU = 0.7 / 0.5) Pedestrian (IoU = 0.5 / 0.25) Cyclist (IoU = 0.5 / 0.25)
@@
Method M Sync. (AP@IoU = 0.7 / 0.5) Async. (AP@IoU = 0.7 / 0.5) AM
@@
Method Modality Vehicle (IoU = 0.5) Pedestrian (IoU = 0.25) Cyclist (IoU = 0.25)
@@ Appendix A Summary
  • Appendix D compares the V2X-Radar single-agent subsets with representative autonomous driving datasets, highlighting their advantages in 4D Radar sensing, adverse-weather coverage, and multi-pass data collection.
  • Appendix E introduces quantitative calibration metrics for LiDAR-Camera and LiDAR-4D Radar pairs, providing reprojection and alignment error analyses that demonstrate the dataset’s high calibration precision.
  • Appendix F showcases visual examples of diverse data collection scenarios, including corner cases for single-vehicle perception and variations across different times of day and weather conditions.
@@ Appendix C Robustness Evaluation of Time Synchronization
To comprehensively evaluate the robustness of existing cooperative perception methods under temporal misalignment, we adopt the time synchronization protocol used in DAIR-V2X [54] and V2V4Real [40]. Specifically, we simulate transmission delays of 0 ms (synchronous), 100 ms, 200 ms, 300 ms, and 400 ms between vehicle-side and roadside sensors on the V2X-Radar-C dataset to analyze the sensitivity of perception performance to temporal offsets. The results, summarized in Tab. 7, demonstrate a consistent degradation trend as the delay increases, highlighting the detrimental impact of temporal asynchrony on cooperative perception.
@@
Appendix D Dataset Comparison from a Single-agent Perspective
@@
Appendix E Quantitative Calibration Metrics
For LiDAR-Camera calibration, the reprojection error (RPE) is used as the evaluation metric. A checkerboard target is employed to extract 3D corner points in the LiDAR frame and their corresponding pixel coordinates in the camera image. Using the estimated calibration parameters, the 3D points are projected onto the image plane, and the mean pixel-wise L2 distance between the projected and true corner locations is reported as the RPE.
For LiDAR-4D Radar calibration, the alignment error (AE) is used. Corner reflectors are employed to obtain paired 3D points in both LiDAR and Radar frames. Radar points are transformed into the LiDAR coordinate system based on the estimated extrinsics, and the mean Euclidean (L2) distance between the transformed Radar points and their corresponding LiDAR points is computed as the AE.
@@ Appendix E Quantitative Calibration Metrics
Appendix F Data Collection Scenarios
Our data collection covers a wide range of weather conditions and times of day, with a particular focus on challenging corner cases for single-vehicle autonomous driving.
Figure 9 further depicts the diversity of the collection scenarios, spanning multiple times of day and varying meteorological conditions As shown in Figure 9(a), the temporal segments include morning, afternoon, dusk, and nighttime scenes. Figure 9(b) demonstrates the range of weather conditions, from clear and sunny to fog, rain, and snow.
@@ Appendix F Data Collection Scenarios
Figure 10 provides a comprehensive visualization of representative corner cases captured in the V2X-Radar, reflecting the diverse and complex challenges encountered in real-world autonomous driving. These include occluded vehicles at public intersections, partially hidden cyclists at T-junctions, concealed pedestrians within restricted park areas, and obscured cyclists at campus crossings.

### v4

@@ 1 Introduction
Perception is critical in autonomous driving. While many single-vehicle perception methods [36; 35; 16; 29; 68; 53; 26; 25; 30; 15; 54; 33; 40; 34; 7; 32; 65] have emerged, substantial safety challenges persist due to occlusions and limited perception ranges. These issues occur because vehicles can only view their surroundings from a single perspective, resulting in an incomplete understanding of the scenario. This limitation hinders autonomous vehicles from achieving safe navigation and making optimal decisions. To address this challenge, recent studies have explored cooperative perception, where ego vehicles extend their perception range and overcome occlusions with assistance from other vehicles or roadside perception [51; 50; 52; 55; 39].
@@
  • We present V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar for cooperative perception. The dataset is collected across diverse real-world scenarios with multiple sensor modalities, comprehensively covering various lighting and weather conditions to enable robust perception evaluation.
@@
  • Comprehensive benchmarks of recent perception algorithms are conducted across cooperative perception on V2X-Radar-C, roadside perception on V2X-Radar-I, and vehicle-side perception on V2X-Radar-V subsets.
@@ 2.1 Autonomous Driving Datasets
Public datasets [1; 31; 48; 47; 14; 62; 60; 46; 41; 20; 66; 67; 37] have significantly accelerated the progress of autonomous driving in recent years. Single-vehicle datasets, such as KITTI [6], NuScenes [1], and Waymo [31], have notably furthered the development of single-vehicle perception. In terms of cooperative perception datasets, OPV2V [48] is the first dataset in this area, gathering data through co-simulation using CARLA [3] and OpenCDA [43; 44]. V2XSet [47] and V2X-Sim [14] explore V2X perception by employing synthesized data from the CARLA simulator [3]. Unlike simulated datasets, DAIR-V2X introduces the first real-world dataset for cooperative detection. V2X-Seq [62] extends sequences from DAIR-V2X [60] with track IDs, establishing a sequential perception and trajectory forecasting dataset. V2V4Real [46] presents the initial real-world V2V dataset gathered from two connected vehicles. V2X-Real [41] is a large-scale multi-modal multi-view dataset intended for V2X research, compiled using two connected automated vehicles and two intelligent roadside units. TUMTraf-V2X [69] is a multimodal, multi-view V2X cooperative perception dataset designed for 3D object detection and tracking in traffic scenarios. A common limitation among the cooperative perception datasets mentioned above is their narrow focus on camera and LiDAR sensors, overlooking the benefits of 4D Radar. 4D Radar is recognized for its superior adaptability to adverse weather conditions, as demonstrated by single-vehicle autonomous driving datasets such as K-Radar [20] and Dual-Radar [66]. Consequently, It’s essential to create a multimodal dataset that integrates 4D Radar to enhance cooperative perception research.
@@ 2.2 Cooperative Perception
Cooperative perception aims to extend the range of perception and overcome occlusions in single-vehicle perception by leveraging shared information among connected agents. It can be classified into three main types based on fusion strategies: (1) Early Fusion, which involves transmitting raw sensor data for the ego vehicle to aggregate and predict objects; (2) Late Fusion, which integrates detection results for a consistent prediction; and (3) Intermediate Fusion, which shares and fuses intermediate high-level features. Recent state-of-the-art methods [38; 2; 45; 47; 9; 18; 38; 4; 61; 63; 5; 28] typically follow intermediate fusion, achieving the best trade-off between accuracy and bandwidth requirements. For instance, V2VNet [38] uses a graph neural network to refine features and perform joint perception and prediction iteratively. F-Cooper [2] introduces a pooling mechanism for identifying salient features. CoBEVT [45] introduced local-global sparse attention for improved cooperative BEV map segmentation. V2X-ViT [47] presented a unified vision transformer for multi-agent and multi-scale perception. HM-ViT [9] introduced a 3D heterogeneous graph transformer for camera and LiDAR fusion. FFNet [61] is a novel flow-based feature fusion framework that addresses temporal asynchrony in cooperative 3D object detection through feature flow prediction. HEAL [18] proposed an extensible collaborative perception framework to integrate new heterogeneous agents with minimal integration cost and performance decline.
@@ 4.2 Cooperative 3D Object Detection
Evaluation Metrics. The evaluation area spans [-100, 100] metres in both the x and y directions relative to the ego vehicle. Similar to DAIR-V2X [60] and V2V4Real [46], we group various vehicle types into the same class and concentrate solely on vehicle detection. The performance of object detection is evaluated using Average Precision (AP) at IoU thresholds of 0.5 and 0.7. Transmission costs are calculated through Average MegaBytes (AM). In line with previous studies [60; 46], we compare methods in two configurations: (1) Synchronous, ignoring communication delays. (2) Asynchronous, simulating the delay by retrieving roadside samples with the preceding timestamp.
@@ 5.2 Benchmark Results
Method M Vehicle (IoU = 0.7 / 0.5) $\uparrow$ Pedestrian (IoU = 0.5 / 0.25) $\uparrow$ Cyclist (IoU = 0.5 / 0.25) $\uparrow$
@@
Method M Vehicle (IoU = 0.7 / 0.5) $\uparrow$ Pedestrian (IoU = 0.5 / 0.25) $\uparrow$ Cyclist (IoU = 0.5 / 0.25) $\uparrow$
@@
Method M Sync. (AP@IoU = 0.7 / 0.5) $\uparrow$ Async. (AP@IoU = 0.7 / 0.5) $\uparrow$ AM
@@
Method Modality Vehicle (IoU = 0.5) $\uparrow$ Pedestrian (IoU = 0.25) $\uparrow$ Cyclist (IoU = 0.25) $\uparrow$
@@ 7 Limitation and Future Work
Acknowledgment
This work was supported by the National Natural Science Foundation of China (52221005, 62273198, 52072215, U1964203), the Beijing Natural Science Foundation (L241017, L243025), the National Key Research and Development Program of China (2022YFB2503003). the Agency for Science, Technology and Research (A*STAR), Singapore, through the MTC Individual Research Grant (M22K2c0079), the CoE Dean’s Interdisciplinary Grant at Nanyang Technological University, and the Ministry of Education, Singapore, through the Tier 2 Grant (MOE-T2EP50222-0002).
@@ Appendix A Summary
  • Appendix D presents the ablation study on the contribution of Doppler velocity information, showing that incorporating Doppler significantly enhances 4D Radar perception under adverse weather conditions.
  • Appendix E compares the V2X-Radar single-agent subsets with representative autonomous driving datasets, highlighting their advantages in 4D Radar sensing, adverse-weather coverage, and multi-pass data collection.
  • Appendix F introduces quantitative calibration metrics for LiDAR-Camera and LiDAR-4D Radar pairs, providing reprojection and alignment error analyses that demonstrate the dataset’s high calibration precision.
  • Appendix G showcases visual examples of diverse data collection scenarios, including corner cases for single-vehicle perception and variations across different times of day and weather conditions.
@@ Appendix C Robustness Evaluation of Time Synchronization
To systematically assess the robustness of existing cooperative perception methods under temporal misalignment, we follow the time synchronization protocol adopted in DAIR-V2X [60] and V2V4Real [46], ensuring methodological consistency with established large-scale benchmarks. Specifically, we simulate varying transmission delays of 0 ms (synchronous), 100 ms, 200 ms, 300 ms, and 400 ms between vehicle-side and roadside sensors on the V2X-Radar-C dataset to comprehensively analyze how communication latency affects perception accuracy. These delay configurations are designed to emulate realistic transmission conditions in vehicular networks. As summarized in Tab. 7, the results show a clear and monotonic degradation trend as the delay increases. This consistent performance decline highlights the detrimental effect of temporal misalignment arising from transmission latency on multi-agent perception, indicating that even moderate desynchronization can substantially disrupt spatial–temporal feature correspondence.
@@
Table 8: Ablation results on Doppler’s contribution to 4D Radar perception under adverse weather (rain, fog, and snow). 4D Radar with Doppler yields consistent gains over the Doppler-free variant (+3.41 AP for vehicles) and even surpasses LiDAR under adverse weather (43.12 vs. 44.80).
[…]
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
LiDAR 47.23 43.12 42.15 21.12 17.52 16.87 29.32 25.34 24.81
4D Radar w/o Doppler 45.22 41.39 40.92 19.98 14.30 14.24 24.68 21.52 20.56
4D Radar w/ Doppler 48.35 44.80 43.91 20.14 16.54 16.20 27.85 23.66 22.68
Appendix D Role of Doppler in 4D Radar Perception
To thoroughly assess the impact of Doppler velocity information, we conduct ablation studies comparing 4D Radar with and without Doppler measurements on a subset of vehicle-side frames captured under adverse weather conditions, including rain, fog, and snow. These challenging scenarios are deliberately chosen to evaluate the robustness of radar-based perception when visual sensors are impaired by low visibility or specular noise. As summarized in Tab. 8, incorporating Doppler cues consistently improves detection performance across all object categories, delivering substantial gains for vehicles (+3.41 AP) and noticeable improvements for pedestrians and cyclists. The performance boost stems from Doppler’s capacity to explicitly encode radial velocity, enabling better motion discrimination and mitigating background clutter interference. Compared with LiDAR, 4D Radar enhanced with Doppler not only narrows the performance gap but even surpasses LiDAR in severe weather (e.g., Vehicle: 43.12 vs. 44.80), underscoring its resilience to optical degradation. Overall, these results demonstrate that Doppler information is not merely an auxiliary signal but a critical sensing dimension that fundamentally enhances the reliability and environmental adaptability of radar perception.
Appendix E Dataset Comparison from a Single-agent Perspective
@@
Appendix F Quantitative Calibration Metrics
For LiDAR-Camera calibration, the reprojection error (RPE) is used as the evaluation metric. A checkerboard target is employed to extract 3D corner points in the LiDAR frame and their corresponding pixel coordinates in the camera image. Using the estimated calibration parameters, the 3D points are projected onto the image plane, and the mean pixel-wise L2 distance between the projected and true corner locations is reported as the reprojection error (RPE).
For LiDAR-4D Radar calibration, the alignment error (AE) is used. Corner reflectors are employed to obtain paired 3D points in both LiDAR and Radar frames. Radar points are transformed into the LiDAR coordinate system based on the estimated extrinsics, and the mean Euclidean (L2) distance between the transformed Radar points and their corresponding LiDAR points is computed as the alignment error (AE).
Appendix G Data Collection Scenarios
Fig. 9 illustrates the diversity of data collection scenarios in the vehicle-side subset, spanning different times of day and a wide range of meteorological conditions. Specifically, Fig. 9(a) depicts temporal variations covering morning, afternoon, dusk, and nighttime scenes, while Fig. 9(b) presents weather diversity, including clear, foggy, rainy, and snowy environments. This comprehensive coverage enables the dataset to capture a broad spectrum of illumination and weather conditions, thereby supporting a more rigorous and comprehensive evaluation of perception robustness under diverse real-world challenges.
@@ Appendix G Data Collection Scenarios
Fig. 10 provides a comprehensive visualization of representative corner cases captured in the V2X-Radar dataset, illustrating the diverse and complex challenges encountered in real-world autonomous driving scenarios. These cases include occluded vehicles at public intersections, partially hidden cyclists at T-junctions, concealed pedestrians within restricted park areas, and obscured cyclists at campus crossings, all of which highlight the need for robust perception under diverse and safety-critical conditions.